OpenStudy (anonymous):

factor y^5+y^4-16y-16

hero (hero):

y^5+y^4-16y-16 y^4(y+1)-16(y+1) (y+1)(y^4-16) (y+1)(y^2+4)(y^2-4) (y+1)(y^2+4)(y+2)(y-2)

hero (hero):

any questions?

OpenStudy (anonymous):

how did you get y+1??

hero (hero):

y^4/y^4 = 1 y^5/y^4 = y^(5-4) = y^1 = y

hero (hero):

-16/-16 = 1 -16y/-16 = y

hero (hero):

When you factor, you're really dividing